OpenStudy (ujjwal):

If y=(x+1)(x+2)(x+3)......(x+n) then dy/dx at x=0 is: a)n! ln(1+x) b)n! ln2 c)n! ln3 d)n! (1+\(\frac{1}{2}\)+\(\frac{1}{3}\)+.... +\(\frac{1}{n}\))

OpenStudy (zarkon):

only one of the answers makes sense

OpenStudy (zarkon):

a) has an x in it...that's not good

OpenStudy (freckles):

You can do it for like small values of n and see if you see one of the patterns above like for n=1,n=2,n=3,n=4 I would stop at n=4 (also hint: don't multiply the junk out after applying product rule)

OpenStudy (zarkon):

b) and c) will give answers that are not integers

OpenStudy (freckles):

Thats true lol

OpenStudy (zarkon):

te derivative of y evaluated at zero will clearly be an integer

OpenStudy (ujjwal):

How do you know that the derivative of y evaluated at zero will be an integer?

OpenStudy (freckles):

Have you done any of the cases I mentioned to convince yourself?

OpenStudy (ujjwal):

I haven't done such problems before..

OpenStudy (freckles):

If y=x+1, then y'=1 so y'(at x=0) =1 If y=(x+1)(x+2), then y'=? <---use product rule....after using product rule don't multiply the junk out and then evaluate y' at x=0 and see what you get

OpenStudy (zarkon):

let n=1

OpenStudy (zarkon):

only one answer will work

OpenStudy (ujjwal):

@Zarkon , you mean to say that we can put value n=1 and see which option satisfies.. And that's it..??

OpenStudy (freckles):

Remember the product of integers will be integers and the sum of integers will be integers

OpenStudy (zarkon):

if they give you options...which they do...yes

OpenStudy (zarkon):

at least for this problem

OpenStudy (ujjwal):

Thanks a lot @Zarkon and @freckles ...